# When Models Don’t Collapse: On the Consistency of Iterative MLE

### v1

@@ 5.1 Models Can Collapse Immediately
In the above theorem, as the number of samples grows, we can find a family of distributions such that $p_{\theta^{(1)}}$ is very close to $p_{\theta^{\star}}$ with high probability, but there is some constant probability that $p_{\theta^{(2)}}$ will be far from $p_{\theta^{(1)}}$. This implies that statements similar to Thm. 4.1 are not possible for general TV-consistent families without further assumptions. Indeed, Thm. 5.1 implies that the relative gap in total variation between iterations $t=1$ and $t=2$ can be arbitrarily large, since
@@
We now provide some intuition for the proof of Thm. 5.1, with the full rigorous proof appearing in Appendix E. We consider a family of distributions, given by the following parameterized mixture of uniform distributions on $\mathbb{R}$:
@@
We now move on to analyzing the second iteration. Because $\alpha^{(1)}\approx\frac{1}{2n}$, then with some constant probability (over the sampling of $n$ new samples from $p_{\theta^{(1)}}$), at least one of these samples $x^{(1)}_{i}$ will be inside the interval $[2,3]$. When this happens, because $f(\alpha)$ is tiny for larger values of $\alpha$ (leading to a high likelihood in the interval $[\mu,\mu+f(\alpha)]$), the MLE solution $\theta^{(2)}=(\alpha^{(2)},\mu^{(2)})$ will be such that $x^{(1)}_{i}\in[\mu^{(2)},\mu^{(2)}+f(\alpha^{(2)})]$ and $\alpha^{(2)}$ will be sufficiently large so that $f(\alpha^{(2)})$ is very small. In particular, $\alpha^{(2)}$ will be considerably larger than the ground truth $\alpha^{(0)}=0$, leading to model collapse.
@@ 5.2 Arbitrarily Fast Model Collapse
Thm. 5.1 shows that without further assumptions, model collapse can occur already after a single iteration. However, the construction requires picking the distribution according to the sample size $n\in\mathbb{N}$, which is arguably unnatural. Below, we show that this requirement can be removed: Namely, there exists a family of distributions where model collapse will occur for any sample size $n$. On the flip side, the model collapse no longer occurs after a single iteration, but rather after a number of iterations that grows with $n$ (although the growth rate can be arbitrarily slow):
@@
Importantly, $\phi$ can be chosen to grow arbitrarily slowly. For example, taking $\phi(n)=\log\log(n+1)$, Thm. 5.2 implies that one can exhibit model collapse in as few as $\mathcal{O}(\log\log(n+1))$ iterations.
The proof of Thm. 5.2 draws inspiration from the proof of Thm. 5.1, but the construction is more involved, as the distribution can no longer depend on the number of samples $n$. The family will consist of two types of distributions. The first, which we denote as $h_{\boldsymbol{\alpha}}$, has the form
@@
Thus, the general procedure is as follows: For any $J\in\mathbb{N}$, once $\alpha_{j}^{(t)}>0$ for every $j\leq J-1$, there is a non-zero chance that a new sample $x^{(t)}_{i}$ will reach interval $[J,J+1]$, ensuring $\alpha_{J}^{(t+1)}>0$. We choose the function $f$ so that for any $N\in\mathbb{N}$, there is some $J_{N}$ such that if $n\leq N$ and if there is some sample in $[J_{N},J_{N}+1]$, then the MLE will be of the form $g_{\beta,J}$ (as $f(J_{N})$ is sufficiently small, leading to a high likelihood of the sample). We show that once this happens, the total variation distance will be large, and the proof will be complete.
The difficult part is showing that for any $J\in\mathbb{N}$, there is some time $T\in\mathbb{N}$ such that with high probability, there will be a sample in $[J,J+1]$. Moreover, this $T$ can be chosen to be essentially independent of $n$. Since we may let $J_{N}$ grow arbitrarily slowly in $N$, and the number of iterations needed to obtain a sample in $[J_{N},J_{N}+1]$ can be upper bounded independently of $N$, the number of iterations needed for model collapse can grow arbitrarily slowly with $N$ (and thus with $n$).
@@ Appendix A Background on Likelihood Estimation
[…]
@@
where the remainder $R(\mathbf{x})$ can be shown to satisfy $|R(\mathbf{x})|\leq\frac{K_{3}}{6}\left\|\theta-\theta^{(0)}\right\|^{3}$ under Assumption  2. By Thm. A.1 the expected value of the gradient term is 00 and by Thm. A.2 the expected value of the hessian term is $-\mathcal{I}(\theta^{(0)})$.
As such, Taylor expanding at every point $\mathbf{x}$ together with Thm. A.1 and Thm. A.2, the KL divergence can be approximated as
@@ Appendix B Concentration
We bring Thm. B.3 to a slightly more convenient form for our uses.
@@
Set $A_{t}^{2}:=K^{2}I_{d}$, $\sigma^{2}:=TK^{2}$, apply Thm. B.2 once to bound $\sum_{t}M_{t}$ and again to bound $-\sum_{t}M_{t}$. The corollary follows from the union bound. ∎
@@
By Thm. A.1, for all $k\in[Tn]$, $\mathbb{E}[\mathbf{z}_{\rho(k)}\mid\mathbf{z}_{\rho(1)},\ldots,\mathbf{z}_{\rho(k-1)}]=\boldsymbol{0}$. Furthermore, by Assumption  2. A, for any $t,i$ and any $u\geq 0$
@@
As such, by Thm. B.1 there exists an absolute constant $C>0$ such that with probability at least $1-\delta$,
@@
We can also obtain concentration for a single $\bar{\theta}\in B_{r}(\theta^{(0)})$. We omit the proof as it is a simplified version of Lemma B.1 (specifically, the assumptions and Thm. A.1 imply that the conditions of Thm. B.1 are satisfied, which gives the following result).
@@
By Thm. A.2, for all $k\in[Tn]$, $\mathbb{E}[M_{\rho(k)}\mid M_{\rho(1)},\ldots,M_{\rho(k-1)}]=\boldsymbol{0}$. Furthermore, by Assumption  2. B and Corollary A.1, for any $k\in[Tn]$,
@@
As such, by Thm. B.3 there exists an absolute constant $C>0$ such that with probability at least $1-\delta$,
@@ C.1 Non-Asymptotic Consistency
[…]
@@
[…]
@@
We now prove the following proposition, which will serve a substantial role in the proof of Thm. 4.1.
@@
For the linear term, first note that if $t\geq 2$ then $\theta^{(t-1)}$ is a stationary point of $\sum_{j=0}^{t-2}\ell_{j}(\cdot)$, so $\sum_{j=0}^{t-2}\nabla\ell_{j}(\theta^{(t-1)})^{\top}=0$. So for any $t\in\mathbb{N}$, $\sum_{j=0}^{t-1}\nabla\ell_{j}(\theta^{(t-1)})=\nabla\ell_{t-1}(\theta^{(t-1)})$. Using this and Lemma B.2, there exists a constant $C_{1}:=C_{1}(K_{1})>0$ such that with probability at least $21-\delta/$ 2,
@@
For the quadratic term, since the matrix $\nabla^{2}\ell_{j}(\theta^{t-1})$ and the Fisher information matrices are symmetric, we have by Weyl’s inequality, Assumption  3 and Lemma C.2 that for $C_{2}=C_{2}(K_{2},K_{3})>0$ it holds with probability at least $21-\delta/$ 2 that
@@
where $\mathbf{z}_{i}\in B$ (and in particular, $\mathbf{z}_{i}\in\Theta$). Let $R_{t}\in\mathbb{R}^{d\times d}$ be the matrix whose coordinates are given by $[R_{t}]_{i,k}:=\frac{1}{2}\sum_{j=0}^{t}\sum_{r=1}^{d}\frac{\partial^{3}}{\partial\theta_{r}\partial\theta_{k}\partial\theta_{i}}\ell_{j}(\mathbf{z}_{i})(\theta^{(t+1)}-\theta^{(t)})_{r}$. Then Eq. (C.2) implies
@@
For $t=0$ the claim is trivial. Now, assume Eq. (15) holds for $t-1$, and we will prove it holds for $t$.
By Eq. (15), for sufficiently large $c$ and the assumption that $n\geq N$, the conditions of Proposition C.1 are satisfied (if Eq. (15) is not $<\frac{\lambda_{0}}{4C_{2}}$ as Proposition C.1 requires, one can replace $c$ by a suitable larger constant that depends on the same parameters), so it implies that with probability at least $61-\delta_{0}/$ 6 (using the union bound),
@@
It remains to bound the terms in Eq. (D). We will first employ an additional probabilistic bound for the gradient terms. By Lemma B.1, with probability at least $21-\delta_{t}/$ 2
@@
Similarly, by Lemma B.2 and the union bound, it holds with probability at least $61-\delta_{0}/$ 6 that
@@
Under the event $A$, from Eq. (18, D, D) and Lemma C.5, it holds for all $\tau\in\{0,\ldots,t-1\}$ that
@@
[…]
@@
where ()($\star$)( ⋆ ) follows whenever $\sqrt{c}\geq C^{\prime}$ by the assumption that
@@
Equivalently, letting $\mathbb{I}$ denote the indicator function (where for any set $A$, $\mathbb{I}_{A}(x)$ is $1$ if $x\in A$ and 00 otherwise), the PDFs $p_{\theta}$ are given by:
@@
Let $x_{\max}:=\max_{i\in[n]}x^{(0)}_{i}$. Note that whenever $\alpha\leq\frac{1-x_{\max}}{2}$, then every $x^{(0)}_{i}$ is inside the interval $[0,1-2\alpha]$. Consequently, for all $\alpha\in\left[0,\frac{1-x_{\max}}{2}\right]$, $\ell_{0}(\theta)=-\log\left(\frac{1}{2}\right)-\log\left(1+\frac{1-\alpha}{1-2\alpha}\right)$. Since the function $-\log\left(1+\frac{1-\alpha}{1-2\alpha}\right)$ is monotonically decreasing in $\alpha$ for all $\alpha<\frac{1}{2}$, $\ell_{0}(\theta)$ is also monotonically decreasing on $\left[0,\frac{1-x_{\max}}{2}\right]$. As such, the MLE $\theta^{(1)}=(\alpha^{(1)},\mu^{(1)})$ which minimizes $\ell_{0}(\theta)$ must satisfy
@@
[…]
@@
Now let $\bar{\alpha}=1/8$ and fix $\bar{\mu}$ such that there is at least one sample in $[\bar{\mu},\bar{\mu}+f(\bar{\alpha})]$ (which we know exists as there is some $x^{(1)}_{i}\in[2,3]$). Plugging this $\bar{\theta}=(\bar{\alpha},\bar{\mu})$ into Eq. (1), using that the first term is negative, and that there is at least one $x^{(1)}_{i}\in[2,3]$, we have:
@@
[…]
@@
By Remark 1, for any $\theta\in\Theta$ and $j\geq b$, the parameters $\alpha_{j}$ do not affect the log likelihood. Furthermore, since $g_{\beta,J}=\frac{1}{2J}\mathbb{I}_{[0,J]}(x)$ the log likelihood is strictly decreasing in $J$ for $\forall J\geq b+1$. As such, for the purpose of showing TV-consistency, we may "discard" all values of $J\geq b+1$ and all indices $\geq J+1$ in $\boldsymbol{\alpha}$, treating $\Theta$ as $[0,\frac{1}{4}]^{J+1}\times[0,1]\times{2,\ldots,J+1}\times\{0,1\}$. This is a closed and bounded subset of a Euclidean space and is therefore compact. Furthermore, $\log p_{\theta^{(0)}}(x)$ are uniformly bounded. So by Lemma F.10, the MLE is consistent. ∎
@@
Now, assume the claim holds for some time $t-1$. Then $\alpha_{j}^{(t)}=\alpha_{j}^{(t_{j}+1)},$ so following Remark 1, all new samples $X^{(t)}$ that are inside the interval $[j,j+1]$ must also be inside the interval
@@
let $B$ denote the event that $\exists i\in[n]\text{ s.t }x^{(t)}_{i}\in[j,j+1]$ and for any $q\in\mathbb{N}$, let $A_{q}$ denote the event that |$\left|\{i\in[n]\mid x^{(t)}_{i}\in[j,j+1]\}\right|\geq q$. Then if $s^{(t)}=0$,
@@
Let $b_{i}$ be $1$ if $x^{(t)}_{i}\in[j,j+1]$ and 00 otherwise. Then conditioned on $\theta^{(1)},\ldots,\theta^{(t)}$, $b_{i}$ are i.i.d. Bernoulli random variables with parameter $u$, so applying Lemma F.7 completes the proof. ∎
@@
For any $t,i$ let $b_{t,i}$ be the bernoulli random variables that take the value of $1$ if $x^{(t)}_{i}\in[j+1,j+2]$ and 00 else. By Remark 1, $b_{t,i}$ are Ber$(u)$ random variables. Let $A_{t}$ denote the event that $\forall b_{t,i}=0$. For any $t\geq t_{j}+1$, $x^{(t)}_{i}$ are i.i.d. when conditioned on $\theta^{(1)},\ldots,\theta^{(t)}$, and when $s^{(t)}=0$, we get
@@
By Lemma F.3, if $M_{t,j+1}>0$ for some $t\in\{t_{j}+1,\ldots,t_{j+1}\}$ and if $s^{(t)},\ldots,s^{(t_{j+1}}=0$ then $M_{t_{j+1},j+1}>0$. In summary, we have given the lower bound on $M_{t_{j+1},j+1}$ needed for the lemma with probability at least $21-\delta/$ 2.
We now move on to the upper bound of the lemma. Suppose that there exists a $\tau$ which is the first timestep for which $M_{\tau,j+1}>0$ or $s^{(\tau)}=1$. If $s^{(\tau)}=1$ we are done, so assume it is 00. Let $B$ denote the event that $\exists i\in[n]\text{ s.t }x^{(\tau)}_{i}\in[j+1,j+2]$ and let $A$ denote the event that |$\left|\{i\in[n]\mid x^{(\tau)}_{i}\in[j+1,j+2]\}\right|\leq q$. We want to bound $M_{\tau,j+1}$, where we must condition on the fact there is at least one sample at time $\tau$ that reached interval $j$. Recall that by Lemma F.3, $\alpha_{k}^{(\tau)}=\alpha_{k}^{(t_{j}+1)}$ for all $k\leq j$. By Lemma F.4, using our choice of $q$ we obtain
@@
By construction, $p_{\theta^{(0)}}$ is the uniform distribution on $[0,1]$, so $M_{0,0}$ is the maximum of $n$ i.i.d. standard uniform random variables. We thus use Lemma F.8 to bound $M_{0,0}$; so it holds with probability at least $21-\delta/$ 2 that
@@
Note that by Lemma F.3, the same bound holds for any $t\geq t_{j}$ such that $s^{(t_{j})},\ldots,s^{(t)}=0$. It was already shown for $j=0$ that $M_{t_{0},0}>0$, so for each $j\in[J]$, conditioning on $t_{0},\ldots,t_{j-1}$ it holds with probability at least $1-j\delta/(2J)$ that either there is some $t\leq t_{j}$ with $s^{(t)}=1$, or the bound on $M_{t_{j},j}$ given in Eq. (32) holds. Applying the union bound, this is true for all $j\in[J]$ with probability at least $21-\delta/$ 2. From now suppose that for all $t\leq t_{J}$, $s^{(t)}=0$ (otherwise we are done).

### v2

@@ 5.1 Models Can Collapse Immediately
In the above theorem, as the number of samples grows, we can find a family of distributions such that $p_{\theta^{(1)}}$ is very close to $p_{\theta^{\star}}$ with high probability, but there is some constant probability that $p_{\theta^{(2)}}$ will be far from $p_{\theta^{(1)}}$. This implies that statements similar to Theorem 4.1 are not possible for general TV-consistent families without further assumptions. Indeed, Theorem 5.1 implies that the relative gap in total variation between iterations $t=1$ and $t=2$ can be arbitrarily large, since
@@
We now provide some intuition for the proof of Theorem 5.1, with the full rigorous proof appearing in Appendix E. We consider a family of distributions, given by the following parameterized mixture of uniform distributions on $\mathbb{R}$:
@@
We now move on to analyzing the second iteration. Because $\alpha^{(1)}\approx\frac{1}{2n}$, then with some constant probability (over the sampling of $n$ new samples from $p_{\theta^{(1)}}$), at least one of these samples $x^{(1)}_{i}$ will be inside the interval $[2,3]$. When this happens, because $f(\alpha)$ is tiny for larger values of $\alpha$ (leading to a high likelihood in the interval $\mu,\mu+f(\alpha)]$), the MLE solution $\theta^{(2)}=(\alpha^{(2)},\mu^{(2)})$ will be such that $x^{(1)}_{i}\in[\mu^{(2)},\mu^{(2)}+f(\alpha^{(2)})]$ and $\alpha^{(2)}$ will be sufficiently large so that $f(\alpha^{(2)})$ is very small. In particular, $\alpha^{(2)}$ will be considerably larger than the ground truth $\alpha^{(0)}=0$, leading to model collapse.
@@ 5.2 Arbitrarily Fast Model Collapse
Theorem 5.1 shows that without further assumptions, model collapse can occur already after a single iteration. However, the construction requires picking the distribution according to the sample size $n\in\mathbb{N}$, which is arguably unnatural. Below, we show that this requirement can be removed: Namely, there exists a family of distributions where model collapse will occur for any sample size $n$. On the flip side, the model collapse no longer occurs after a single iteration, but rather after a number of iterations that grows with $n$ (although the growth rate can be arbitrarily slow):
@@
Importantly, $\phi$ can be chosen to grow arbitrarily slowly. For example, taking $\phi(n)=\log\log(n+1)$, Theorem 5.2 implies that one can exhibit model collapse in as few as $\mathcal{O}(\log\log(n+1))$ iterations.
The proof of Theorem 5.2 draws inspiration from the proof of Theorem 5.1, but the construction is more involved, as the distribution can no longer depend on the number of samples $n$. The family will consist of two types of distributions. The first, which we denote as $h_{\boldsymbol{\alpha}}$, has the form
@@
Thus, the general procedure is as follows: For any $J\in\mathbb{N}$, once $\alpha_{j}^{(t)}>0$ for every $j\leq J-1$, there is a non-zero chance that a new sample $x^{(t)}_{i}$ will reach interval $[J,J+1]$, ensuring $\alpha_{J}^{(t+1)}>0$. We choose the function $f$ so that for any $N\in\mathbb{N}$, there is some $J_{N}$ such that if $n\leq N$ and if there is some sample in $J_{N},J_{N}+1]$, then the MLE will be of the form $g_{\beta,J}$ (as $f(J_{N})$ is sufficiently small, leading to a high likelihood of the sample). We show that once this happens, the total variation distance will be large, and the proof will be complete.
The difficult part is showing that for any $J\in\mathbb{N}$, there is some time $T\in\mathbb{N}$ such that with high probability, there will be a sample in $[J,J+1]$. Moreover, this $T$ can be chosen to be essentially independent of $n$. Since we may let $J_{N}$ grow arbitrarily slowly in $N$, and the number of iterations needed to obtain a sample in $J_{N},J_{N}+1]$ can be upper bounded independently of $N$, the number of iterations needed for model collapse can grow arbitrarily slowly with $N$ (and thus with $n$).
5.3 Implications and Relation to Theorem 4.1
The results of this section inform us how in the absence of the assumptions of Theorem 4.1, model collapse can occur arbitrarily quickly. Even though the constructions of Theorems 5.1 and 5.2 are artificial, they highlight more general phenomena that are needed for model collapse to occur or to be avoided. In particular, in our view, the main difference from Theorem 4.1 is the smoothness assumption. Theorems 5.1 and 5.2 crucially use a highly non-smooth construction, in which slight perturbations of the parameters can induce huge differences in the resulting model, and we find it unlikely that a negative example would be possible without this behavior.
@@ Appendix A Background on Likelihood Estimation
$\operatorname{TV}\left(p,q\right):=\frac{1}{2}\int_{x\in\mathcal{X}}\left|p(x)-q(x)\right|dx.$
@@
where the remainder $R(\mathbf{x})$ can be shown to satisfy $|R(\mathbf{x})|\leq\frac{K_{3}}{6}\left\|\theta-\theta^{(0)}\right\|^{3}$ under Assumption  2. By Theorem A.1 the expected value of the gradient term is 0 and by Theorem A.2 the expected value of the hessian term is $\mathcal{I}(\theta^{(0)})$.
As such, Taylor expanding at every point $\mathbf{x}$ together with Theorem A.1 and Theorem A.2, the KL divergence can be approximated as
@@ Appendix B Concentration
We bring Theorem B.3 to a slightly more convenient form for our uses.
@@
Set $A_{t}^{2}:=K^{2}I_{d}$, $\sigma^{2}:=TK^{2}$, apply Theorem B.2 once to bound $\sum_{t}M_{t}$ and again to bound $\sum_{t}M_{t}$. The corollary follows from the union bound. ∎
@@
By Theorem A.1, for all $k\in[Tn]$, $\mathbb{E}[\mathbf{z}_{\rho(k)}\mid\mathbf{z}_{\rho(1)},\ldots,\mathbf{z}_{\rho(k-1)}]=\boldsymbol{0}$. Furthermore, by Assumption  2. A, for any $t,i$ and any $u\geq 0$
@@
As such, by Theorem B.1 there exists an absolute constant $C>0$ such that with probability at least $1-\delta$,
@@
We can also obtain concentration for a single $\bar{\theta}\in B_{r}(\theta^{(0)})$. We omit the proof as it is a simplified version of Lemma B.1 (specifically, the assumptions and Theorem A.1 imply that the conditions of Theorem B.1 are satisfied, which gives the following result).
@@
By Theorem A.2, for all $k\in[Tn]$, $\mathbb{E}[M_{\rho(k)}\mid M_{\rho(1)},\ldots,M_{\rho(k-1)}]=\boldsymbol{0}$. Furthermore, by Assumption  2. B and Corollary A.1, for any $k\in[Tn]$,
@@
As such, by Theorem B.3 there exists an absolute constant $C>0$ such that with probability at least $1-\delta$,
@@ C.1 Non-Asymptotic Consistency
$\nabla_{\theta}^{2}\log p_{\theta^{\prime}}(\mathbf{x})-\nabla_{\theta}^{2}\log p_{\theta}(\mathbf{x})=\int_{0}^{1}\frac{d}{dt}\nabla_{\theta}^{2}\log p_{\gamma(t)}(\mathbf{x})dt=\int_{0}^{1}\nabla_{\theta}^{3}\log p_{\gamma(t)}(\mathbf{x})[\theta^{\prime}-\theta]dt,$
@@
$\left\|\nabla_{\theta}^{2}\log p_{\theta^{\prime}}(\mathbf{x})-\nabla_{\theta}^{2}\log p_{\theta}(\mathbf{x})\right\|\leq$ $\int_{0}^{1}\left\|\nabla_{\theta}^{3}\log p_{\gamma(t)}(\mathbf{x})[\theta^{\prime}-\theta]\right\|dt$
@@
We now prove the following proposition, which will serve a substantial role in the proof of Theorem 4.1.
@@
For the linear term, first note that if $t\geq 2$ then $\theta^{(t-1)}$ is a stationary point of $\sum_{j=0}^{t-2}\ell_{j}(\cdot)$, so $\sum_{j=0}^{t-2}\nabla\ell_{j}(\theta^{(t-1)})^{\top}=0$. So for any $t\in\mathbb{N}$, $\sum_{j=0}^{t-1}\nabla\ell_{j}(\theta^{(t-1)})=\nabla\ell_{t-1}(\theta^{(t-1)})$. Using this and Lemma B.2, there exists a constant $C_{1}:=C_{1}(K_{1})>0$ such that with probability at least $1-\delta/2$,
@@
For the quadratic term, since the matrix $\nabla^{2}\ell_{j}(\theta^{t-1})$ and the Fisher information matrices are symmetric, we have by Weyl’s inequality, Assumption  3 and Lemma C.2 that for $C_{2}=C_{2}(K_{2},K_{3})>0$ it holds with probability at least $1-\delta/2$ that
@@
where $\mathbf{z}_{i}\in B$ (and in particular, $\mathbf{z}_{i}\in\Theta$). Let $R_{t}\in\mathbb{R}^{d\times d}$ be the matrix whose coordinates are given by $R_{t}]_{i,k}:=\frac{1}{2}\sum_{j=0}^{t}\sum_{r=1}^{d}\frac{\partial^{3}}{\partial\theta_{r}\partial\theta_{k}\partial\theta_{i}}\ell_{j}(\mathbf{z}_{i})(\theta^{(t+1)}-\theta^{(t)})_{r}$. Then Eq. (C.2) implies
@@
For $t=0$ the claim is trivial. Now, assume Eq. (15) holds for t1t-1, and we will prove it holds for $t$.
By Eq. (15), for sufficiently large $c$ and the assumption that $n\geq N$, the conditions of Proposition C.1 are satisfied (if Eq. (15) is not $\frac{\lambda_{0}}{4C_{2}}$ as Proposition C.1 requires, one can replace $c$ by a suitable larger constant that depends on the same parameters), so it implies that with probability at least $1-\delta_{0}/6$ (using the union bound),
@@
It remains to bound the terms in Eq. (D). We will first employ an additional probabilistic bound for the gradient terms. By Lemma B.1, with probability at least $1-\delta_{t}/2$
@@
Similarly, by Lemma B.2 and the union bound, it holds with probability at least $1-\delta_{0}/6$ that
@@
Under the event AA, from Eq. (18, D, D) and Lemma C.5, it holds for all $\tau\in\{0,\ldots,t-1\}$ that
@@
Combining Eq. (24), Eq. (D) and the fact that $\sum_{\tau=1}^{t}\frac{1}{\tau}\leq 1+\int_{1}^{t}\frac{1}{x}dx\leq 1+\log(t)$, we have for a suitable $C^{\prime}=C^{\prime}(K_{1},K_{2},K_{3},\lambda_{0},r)$
@@
where $(\star)$ follows whenever $\sqrt{c}\geq C^{\prime}$ by the assumption that
@@
Equivalently, letting $\mathbb{I}$ denote the indicator function (where for any set $A$, $\mathbb{I}_{A}(x)$ is $1$ if $x\in A$ and 0 otherwise), the PDFs $p_{\theta}$ are given by:
@@
Let $x_{\max}:=\max_{i\in[n]}x^{(0)}_{i}$. Note that whenever $\alpha\leq\frac{1-x_{\max}}{2}$, then every $x^{(0)}_{i}$ is inside the interval $0,1-2\alpha]$. Consequently, for all $\alpha\in\left[0,\frac{1-x_{\max}}{2}\right]$, $\ell_{0}(\theta)=-\log\left(\frac{1}{2}\right)-\log\left(1+\frac{1-\alpha}{1-2\alpha}\right)$. Since the function $\log\left(1+\frac{1-\alpha}{1-2\alpha}\right)$ is monotonically decreasing in $\alpha$ for all $\alpha<\frac{1}{2}$, $\ell_{0}(\theta)$ is also monotonically decreasing on $\left[0,\frac{1-x_{\max}}{2}\right]$. As such, the MLE $\theta^{(1)}=(\alpha^{(1)},\mu^{(1)})$ which minimizes $\ell_{0}(\theta)$ must satisfy
@@
$\operatorname{TV}\left(p_{\theta^{(0)}},p_{\theta^{(1)}}\right)=$ $\frac{1}{2}\int_{0}^{1}\left|1-p_{\theta^{(1)}}\right|(x)dx+\frac{1}{2}\int_{2}^{3}p_{\theta^{(1)}}(x)dx$
$=$ $\frac{1}{2}\int_{0}^{1-2\alpha^{(1)}}\left|1-\left(\frac{1}{2}+\frac{1-\alpha^{(1)}}{2(1-2\alpha^{(1)})}\right)\right|dx+\frac{1}{2}\int_{1-2\alpha^{(1)}}^{1}\left|1-\frac{1}{2}\right|dx+\frac{\alpha^{(1)}}{4}$
@@
Now let $\bar{\alpha}=1/8$ and fix $\bar{\mu}$ such that there is at least one sample in $\bar{\mu},\bar{\mu}+f(\bar{\alpha})]$ (which we know exists as there is some $x^{(1)}_{i}\in[2,3]$). Plugging this $\bar{\theta}=(\bar{\alpha},\bar{\mu})$ into Eq. (1), using that the first term is negative, and that there is at least one $x^{(1)}_{i}\in[2,3]$, we have:
@@
To see that this is a valid PDF, first note that $\int_{-\infty}^{\infty}h_{\boldsymbol{\alpha}}(x)dx=\sum_{j=0}^{\infty}(1-\alpha_{j})\left(\prod_{k=0}^{j-1}\alpha_{k}\right)$. Now consider any fixed $M\in\mathbb{N}$, then
@@
By Remark 1, for any $\theta\in\Theta$ and $j\geq b$, the parameters $\alpha_{j}$ do not affect the log likelihood. Furthermore, since $g_{\beta,J}=\frac{1}{2J}\mathbb{I}_{[0,J]}(x)$ the log likelihood is strictly decreasing in $J$ for $\forall J\geq b+1$. As such, for the purpose of showing TV-consistency, we may "discard" all values of $J\geq b+1$ and all indices $\geq J+1$ in $\boldsymbol{\alpha}$, treating $\Theta$ as $0,\frac{1}{4}]^{J+1}\times[0,1]\times{2,\ldots,J+1}\times\{0,1\}$. This is a closed and bounded subset of a Euclidean space and is therefore compact. Furthermore, $\log p_{\theta^{(0)}}(x)$ are uniformly bounded. So by Lemma F.10, the MLE is consistent. ∎
@@
Now, assume the claim holds for some time t1t-1. Then $\alpha_{j}^{(t)}=\alpha_{j}^{(t_{j}+1)},$ so following Remark 1, all new samples $X^{(t)}$ that are inside the interval $[j,j+1]$ must also be inside the interval
@@
let $B$ denote the event that $\exists i\in[n]\text{ s.t }x^{(t)}_{i}\in[j,j+1]$ and for any $q\in\mathbb{N}$, let $A_{q}$ denote the event that $\left|\{i\in[n]\mid x^{(t)}_{i}\in[j,j+1]\}\right|\geq q$. Then if $s^{(t)}=0$,
@@
Let $b_{i}$ be $1$ if $x^{(t)}_{i}\in[j,j+1]$ and 0 otherwise. Then conditioned on $\theta^{(1)},\ldots,\theta^{(t)}$, $b_{i}$ are i.i.d. Bernoulli random variables with parameter $u$, so applying Lemma F.7 completes the proof. ∎
@@
For any $t,i$ let $b_{t,i}$ be the bernoulli random variables that take the value of $1$ if $x^{(t)}_{i}\in[j+1,j+2]$ and 0 else. By Remark 1, $b_{t,i}$ are Ber(u)(u) random variables. Let $A_{t}$ denote the event that $\forall b_{t,i}=0$. For any $t\geq t_{j}+1$, $x^{(t)}_{i}$ are i.i.d. when conditioned on $\theta^{(1)},\ldots,\theta^{(t)}$, and when $s^{(t)}=0$, we get
@@
By Lemma F.3, if $M_{t,j+1}>0$ for some $t\in\{t_{j}+1,\ldots,t_{j+1}\}$ and if $s^{(t)},\ldots,s^{(t_{j+1}}=0$ then $M_{t_{j+1},j+1}>0$. In summary, we have given the lower bound on $M_{t_{j+1},j+1}$ needed for the lemma with probability at least $1-\delta/2$.
We now move on to the upper bound of the lemma. Suppose that there exists a $\tau$ which is the first timestep for which $M_{\tau,j+1}>0$ or $s^{(\tau)}=1$. If $s^{(\tau)}=1$ we are done, so assume it is 0. Let $B$ denote the event that $\exists i\in[n]\text{ s.t }x^{(\tau)}_{i}\in[j+1,j+2]$ and let $A$ denote the event that $\left|\{i\in[n]\mid x^{(\tau)}_{i}\in[j+1,j+2]\}\right|\leq q$. We want to bound $M_{\tau,j+1}$, where we must condition on the fact there is at least one sample at time $\tau$ that reached interval $j$. Recall that by Lemma F.3, $\alpha_{k}^{(\tau)}=\alpha_{k}^{(t_{j}+1)}$ for all $k\leq j$. By Lemma F.4, using our choice of $q$ we obtain
@@
By construction, $p_{\theta^{(0)}}$ is the uniform distribution on $[0,1]$, so $M_{0,0}$ is the maximum of $n$ i.i.d. standard uniform random variables. We thus use Lemma F.8 to bound $M_{0,0}$; so it holds with probability at least $1-\delta/2$ that
@@
Note that by Lemma F.3, the same bound holds for any $t\geq t_{j}$ such that $s^{(t_{j})},\ldots,s^{(t)}=0$. It was already shown for $j=0$ that $M_{t_{0},0}>0$, so for each $j\in[J]$, conditioning on $t_{0},\ldots,t_{j-1}$ it holds with probability at least $1-j\delta/(2J)$ that either there is some $t\leq t_{j}$ with $s^{(t)}=1$, or the bound on $M_{t_{j},j}$ given in Eq. (32) holds. Applying the union bound, this is true for all $j\in[J]$ with probability at least $1-\delta/2$. From now suppose that for all $t\leq t_{J}$, $s^{(t)}=0$ (otherwise we are done).